# AnyUp: Universal Feature Upsampling

### v1

@@ 1 Introduction
Bilateral Filtering (Tomasi & Manduchi, 1998)
@@
General image feature extractors, such as DINO (Caron et al., 2021; Oquab et al., 2023; Siméoni et al., 2025), CLIP (Radford et al., 2021), SigLIP (Tschannen et al., 2025), or MAE (He et al., 2022), have become fundamental building blocks in modern day computer vision, providing, e.g., semantics or language-alignment to a wide range of down-stream applications, such as depth estimation or 3D reconstruction (Yang et al., 2024b; Wang et al., 2025), open-vocabulary semantic segmentation (Engelmann et al., 2024; Wysoczańska et al., 2024), or benchmarking generative models (Asim et al., 2025). An important limitation of such pre-trained models, which are usually transformer-based, is that their output feature map resolution is limited to the number of transformer tokens, preventing the prediction of pixel-level features. Therefore, several recent works, such as FeatUp (Fu et al., 2024), LoftUp (Huang et al., 2025), or JAFAR (Couairon et al., 2025) propose learned feature upsampling methods.
While such feature upsampling methods perform well when paired with the vision encoders with which they were trained, they are generally not encoder-agnostic at inference time and need to be retrained to be usable with a different feature extractor. This can be costly or, in the case of the latest large vision models (Siméoni et al., 2025), even infeasible with limited computing resources, as, during training of the upsampler, the vision encoder must be queried at least once on a higher-resolution image for every training sample. In this work, we alleviate this limitation and propose AnyUp, a learned feature upsampling method that generalizes to features of any size provided in any resolution while achieving state-of-the-art performance across various downstream tasks. As shown in Tab. 1, AnyUp is the first of its kind.
@@
In our experiments, AnyUp establishes a new state of the art for feature upsampling. Its key advantages include universal applicability – it can be trained and applied across all feature types and resolutions – and robust generalization to feature types it is was not trained on. Moreover, AnyUp ensures high fidelity by minimizing the distortion of original feature semantics. Through ablations, we analyze different design choices and draw general insights for the feature upsampling task.
@@ 2 Related Work
Learnable upsampling methods (Suri et al., 2024; Fu et al., 2024; Huang et al., 2025; Couairon et al., 2025) try to improve the quality and robustness over learning-free approaches. The parameterization, however, always results in the loss of feature independence at inference time, i.e., upsampling models are fitted to a specific feature encoder (and input dimensionality). In addition, some methods are fixed to a specific downstream task or model (Wang et al., 2019; Lu et al., 2022; Liu et al., 2023; Zhou et al., 2024) or cannot be used with arbitrary scaling factors (Odena et al., 2016; Wang et al., 2019; Lu et al., 2022; Liu et al., 2023; Suri et al., 2024; Fu et al., 2024; Ranzinger et al., 2025). The current state-of-the-art approaches, JAFAR (Couairon et al., 2025) and LoftUp (Huang et al., 2025), both concurrent works, formulate feature upsampling as a single or stacked high-res to low-res attention, respectively. We adopt this formulation as it offers the advantage of being naturally resolution-agnostic, i.e., features can be upsampled from any to any resolution. However, in our AnyUp, we aim to make this architecture encoder-agnostic, i.e., the upsampling model should be trained once and subsequently be applicable to features from any vision encoder, potentially of different feature dimensionality. More details on selected related work can be found in App. A.
@@ 4.2 Local Window Attention
Analyzing the attention patterns of JAFAR (Couairon et al., 2025), we find that in the global attention mechanism, where a pixel query can attend to any feature patch in the input feature map, sometimes vastly unrelated and distant image areas are used as references for upsampling. We argue that we can avoid such patterns and simplify the upsampling problem by restricting the attention computation to local windows around the query point (Ramachandran et al., 2019)., As a high-resolution feature is now a linear combination of a much smaller set of coarse features compared with the global attention, the optimization objective for the upsampler gets easier, as well as efficiency is improved. We refer the reader to App. D for a visualization of the mentioned attention outliers.
@@ 4.3.2 Objective Function
We follow prior work in feature denoising and feature upsampling (Yang et al., 2024a; Couairon et al., 2025) and minimize the distance between predicted features $q^{\prime}$ and target features $\hat{q}$:
@@
In addition, we add a self-consistency regularization $L_{\text{self-consistency}}$, which we detail in App. B.1, as well as an input-consistency regularization $L_{\text{input-consistency}}$, where we simply compute $L_{\text{cos-mse}}$ with the input features $p$ and the accordingly downsampled predicted output features $q$. We do this to further improve the locality of the upsampled features, which is important for sub-object-level tasks like surface normal estimation, as well as maintaining the input feature space.
@@ 5 Experiments
We train AnyUp on the ImageNet dataset (Krizhevsky et al., 2017) and provide further implementation details in App. B. Comparisons against previous works are provided in Sec. 5.1, where we evaluate probing quality, feature space preservation, and behaviour with different resolution changes. Sec. 5.2 provides more results on the generalization of AnyUp to other features at test time. Sec. 5.3 wraps up the experiments with an ablation study.
@@ 5.1 Comparison to Prior Art
Figure 5: Qualitative Probing Results. Visualization of linear probing results for monocular depth estimation on NYUv2 and semantic segmentation on ADE20k. More visualizations given in App. D.
@@
We follow the evaluation scheme proposed by prior works (Couairon et al., 2025, Sec. 4.3.1), where a linear head, i.e., a 1x1 convolution layer, is trained to predict semantic segmentation labels from the high-resolution feature maps. The input images are resized to 448x448 resolution and features are thus of resolution 28x28 or 32x32 depending on the patch size of the vision transformer used for feature extraction. Upsampling is generally performed back to the input image resolution, i.e., equaling a 14x or 16x upsampling, if not mentioned otherwise.
@@
We evaluate performance on the COCO-Stuff (Lin et al., 2014), ADE20k (Zhou et al., 2018) and PASCAL VOC (Everingham et al., 2014) datasets. Results are reported as mean Intersection-over-Union (mIoU) and pixel-wise accuracy. As observed in Tab. 2, AnyUp provides state-of-the-art upsampling performance for downstream semantic segmentation.
@@
Table 3: Surface Normal and Monocular Depth Estimation.
@@
Ideally, the upsampled features should stay in the same space and distribution as the original, low-resolution features. If that is the case, a linear probe trained on the original feature space should directly transfer to the high-resolution features outputted by a given upsampler for the same input feature extractor, without any finetuning. We test how much different feature upsampling methods preserve the feature distribution, by using linear probes pre-trained on the original DINOv2 (ViT-S) feature extractor (Oquab et al., 2023) [^2] We note that pre-trained linear probes and other feature upsampling methods were not released for the more recent DINOv3 (Siméoni et al., 2025) method, hence we stick to DINOv2.. In Tab. 5, we evaluate this for two different tasks, depth estimation on the NYUv2 dataset (Silberman et al., 2012) and semantic segmentation on the ADE20k dataset (Zhou et al., 2018). Qualitative results are in App. D.
AnyUp preserves the input features the best, while improving the prediction quality at higher resolution through improved upsampling, when compared to learnable and heuristic methods. On the other end of the spectrum, predictions using LoftUp are heavily degraded, which can be explained by the affinity matrix loss employed in its training, aligning intra-image feature similarities instead of directly supervising the predictions with the target features (Huang et al., 2025, Sec. 4.2).
@@ 5.2 Feature-Agnostic Upsampling
In Fig. 1, we visualize PCA projections of features, where the same feature upsampling model instance, i.e., a model that was trained only on DINOv2 ViT-S features, is applied on features extracted from image encoders ranging from ResNet (He et al., 2016) to DINOv3 (Siméoni et al., 2025). In Tab. 6, we showcase the remarkable generalization capabilities of AnyUp, where a model trained on DINOv2 and tested on SigLIP 2 outperforms or comes close to the performance of models directly trained on SigLIP 2. Also, a DINOv2-trained AnyUp model generalizes well to DINOv3 features. Finally, as detailed in Fig. 6, AnyUp generalizes well across different DINOv2 architectures.
@@ 6 Limitations
FeatSharp (Ranzinger et al., 2025) combines feature upsampling with explicitly de-noising the input features from positional encoding artifacts (Yang et al., 2024a). We note that this learned per-model denoising step on the input features can be easily prepended to our pipeline but as the learned denoising weights have not been made public at the time of submission, we do not include this analysis in our experiments. Further, through our training strategy on image parts, our model is effectively also trained to mostly ignore such positional encoding artifacts.
@@ 7 Conclusion
We introduced AnyUp, a method for feature upsampling from any resolution to any resolution, which generalizes to feature representations that it was not trained on. Key technical novelties include a feature-agnostic layer, windowed attention, and a training strategy, which work together to achieve state-of-the-art upsampling quality. We make our code and models publicly available at https://github.com/wimmerth/anyup.
@@ Appendix A Additional Details on Related Work
Appendix B Implementation Details
We train our method on the ImageNet dataset (Krizhevsky et al., 2017) for a total of 100,000 training steps. We choose a batch size of 4 with 4 random local crops per training image. Training takes around 5 hours on a single NVIDIA-H100 GPU. We use the AdamW optimizer (Loshchilov & Hutter, 2019) with a learning rate of 2e-4 and a batch size of 4.

### v2

@@ 1 Introduction
Bilateral Filtering (Tomasi and Manduchi, 1998)
@@
General image feature extractors, such as DINO (Caron et al., 2021; Oquab et al., 2023; Xue et al., 2025), CLIP (Radford et al., 2021), SigLIP (Tschannen et al., 2025), or MAE (He et al., 2022), have become fundamental building blocks in modern day computer vision, providing, e.g., semantics or language-alignment to a wide range of down-stream applications, such as depth estimation or 3D reconstruction (Feng et al., 2024; Wang et al., 2025), open-vocabulary semantic segmentation (Engelmann et al., 2024; Wysoczańska et al., 2024), or benchmarking generative models (Asim et al., 2025). An important limitation of such pre-trained models, which are usually transformer-based, is that their output feature map resolution is limited to the number of transformer tokens, preventing the prediction of pixel-level features. Therefore, several recent works, such as FeatUp (Fu et al., 2024), LoftUp (Huang et al., 2025), or JAFAR (Couairon et al., 2025) propose learned feature upsampling methods.
While such feature upsampling methods perform well when paired with the vision encoders with which they were trained, they are generally not encoder-agnostic at inference time and need to be retrained to be usable with a different feature extractor. This can be costly or, in the case of the latest large vision models (Xue et al., 2025), even infeasible with limited computing resources, as, during training of the upsampler, the vision encoder must be queried multiple times for every training sample, often also on a higher-resolution image. In this work, we alleviate this limitation and propose AnyUp, a learned feature upsampling method that generalizes to features of any size provided in any resolution while achieving state-of-the-art performance across various downstream tasks. As shown in Tab. 1, AnyUp is the first of its kind.
@@
In our experiments, AnyUp establishes a new state of the art for feature upsampling. Its key advantages include universal applicability – it can be trained and applied across all feature types and resolutions – and robust generalization to feature types it was not trained on. Moreover, AnyUp ensures high fidelity by minimizing the distortion of original feature semantics. Through ablations, we analyze different design choices and draw general insights for the feature upsampling task.
@@ 2 Related Work
Learnable upsampling methods (Suri et al., 2024; Fu et al., 2024; Huang et al., 2025; Couairon et al., 2025) try to improve the quality and robustness over learning-free approaches. The parameterization, however, always results in the loss of feature independence at inference time, i.e., upsampling models are fitted to a specific feature encoder (and input dimensionality). In addition, some methods are fixed to a specific downstream task or model (Wang et al., 2019; Lu et al., 2022; Liu et al., 2023; Zhou et al., 2025) or cannot be used with arbitrary scaling factors (Odena et al., 2016; Wang et al., 2019; Lu et al., 2022; Liu et al., 2023; Suri et al., 2024; Fu et al., 2024; Ranzinger et al., 2025). The current state-of-the-art approaches, JAFAR (Couairon et al., 2025) and LoftUp (Huang et al., 2025), both concurrent works, formulate feature upsampling as a single or stacked high-res to low-res attention, respectively. We adopt this formulation as it offers the advantage of being naturally resolution-agnostic, i.e., features can be upsampled from any to any resolution. However, in our AnyUp, we aim to make this architecture encoder-agnostic, i.e., the upsampling model should be trained once and subsequently be applicable to features from any vision encoder, potentially of different feature dimensionality. More details on selected related work can be found in App. A.
@@ 4.2 Local Window Attention
Analyzing the attention patterns of JAFAR (Couairon et al., 2025), we find that in the global attention mechanism, where a pixel query can attend to any feature patch in the input feature map, sometimes vastly unrelated and distant image areas are used as references for upsampling. We argue that we can avoid such patterns and simplify the upsampling problem by restricting the attention computation to local windows around the query point (Ramachandran et al., 2019). As a high-resolution feature is now a linear combination of a much smaller set of coarse features compared with the global attention, the optimization objective for the upsampler gets easier, as well as efficiency is improved. We refer the reader to App. E for a visualization of the mentioned attention outliers.
@@ 4.3.2 Objective Function
We follow prior work in feature denoising and feature upsampling (Yang et al., 2024; Couairon et al., 2025) and minimize the distance between predicted features $q^{\prime}$ and target features $\hat{q}$:
@@
In addition, we add a self-consistency regularization $L_{\text{self-consistency}}$, which we detail in App. C.1, as well as an input-consistency regularization $L_{\text{input-consistency}}$, where we simply compute $L_{\text{cos-mse}}$ with the input features $p$ and the accordingly downsampled predicted output features $q$. We do this to further improve the locality of the upsampled features, which is important for sub-object-level tasks like surface normal estimation, as well as maintaining the input feature space.
@@ 5 Experiments
We train AnyUp on the ImageNet dataset (Krizhevsky et al., 2017) and provide further implementation details in App. C. Comparisons against previous works are provided in Sec. 5.1, where we evaluate probing quality, feature space preservation, and behaviour with different resolution changes. Sec. 5.2 provides more results on the generalization of AnyUp to other features at test time. Sec. 5.3 wraps up the experiments with an ablation study.
@@ 5.1 Comparison to Prior Art
Figure 5: Qualitative Probing Results. Visualization of linear probing results for monocular depth estimation on NYUv2 and semantic segmentation on ADE20k. More visualizations given in App. E.
@@
We follow the evaluation scheme proposed by prior works (Couairon et al., 2025, Sec. 4.3.1), where a linear probe, i.e., a 1x1 convolution layer, is trained to predict semantic segmentation labels from the high-resolution feature maps. The input images are resized to 448x448 resolution and features are thus of resolution 28x28 or 32x32 depending on the patch size of the vision transformer used for feature extraction. Upsampling is generally performed back to the input image resolution, i.e., equaling a 14x or 16x upsampling, if not mentioned otherwise. We evaluate performance on the COCO-Stuff (Lin et al., 2014), ADE20k (Zhou et al., 2018) and PASCAL VOC (Everingham et al., 2014) datasets.
@@
Results are reported as mean Intersection-over-Union (mIoU) and pixel-wise accuracy. As observed in Tab. 2, AnyUp provides state-of-the-art upsampling performance for downstream semantic segmentation.
@@
Table 3: Surface Normal and Monocular Depth Estimation. AnyUp outperforms previous upsamplers for geometric tasks. Evaluation on the NYUv2 dataset (Silberman et al., 2012).
@@
Ideally, the upsampled features should stay in the same space and distribution as the original, low-resolution features. If that is the case, a linear probe trained on the original feature space should directly transfer to the high-resolution features outputted by a given upsampler for the same input feature extractor, without any finetuning. We test how much different feature upsampling methods preserve the feature distribution, by using linear probes pre-trained on the original DINOv2 (ViT-S) feature extractor (Oquab et al., 2023) [^2] We note that pre-trained linear probes and other feature upsampling methods were not released for the more recent DINOv3 (Xue et al., 2025) method, hence we stick to DINOv2.. In Tab. 5, we evaluate this for two different tasks, depth estimation on the NYUv2 dataset (Silberman et al., 2012) and semantic segmentation on the ADE20k dataset (Zhou et al., 2018). Qualitative results are in App. E.
AnyUp preserves the input features the best, while improving the prediction quality at higher resolution through improved upsampling, when compared to learnable and heuristic methods. On the other end of the spectrum, predictions using LoftUp are heavily degraded, which can be explained by the affinity matrix loss employed in its training, aligning intra-image feature similarities instead of directly supervising the predictions with the target features (Huang et al., 2025, Sec. 4.2). We include the implicit feature upsampling version of FeatUp (Fu et al., 2024), whose computational cost also prohibits upsampler-specific probe training, in a similar experiment on images with lower resolution in App. B.2.
Computational Performance.
The proposed window-based attention used for the main upsampling operation can be implemented efficiently, reducing the runtime and memory requirements by more than 50% compared to prior works JAFAR and LoftUp, while being slightly less efficient than FeatUp. More details on parameter count, FLOPs and runtime, as well as memory usage in forward and backward passes can be found in App. B.1.
@@ 5.2 Feature-Agnostic Upsampling
In Fig. 1, we visualize PCA projections of features, where the same feature upsampling model instance, i.e., a model that was trained only on DINOv2 ViT-S features, is applied on features extracted from image encoders ranging from ResNet (He et al., 2016) to DINOv3 (Xue et al., 2025). In Tab. 6, we showcase the remarkable generalization capabilities of AnyUp, where a model trained on DINOv2 and tested on SigLIP 2 outperforms or comes close to the performance of models directly trained on SigLIP 2. Also, a DINOv2-trained AnyUp model generalizes well to DINOv3 features. Finally, as detailed in Fig. 6, AnyUp generalizes well across different DINOv2 architectures.
Training on Multiple Feature Extractors.
Our feature-agnostic upsampler architecture allows for training on multiple vision backbones. We make use of this property in an experiment to test whether such training can increase generalization performance to completely unseen features at test time. In this experiment, we train AnyUp on features from the following models: DINOv2 (ViT-S), CLIP (ViT-B), SigLIP (ViT-B), DINOv2 w/ registers (ViT-S), and ViT-B trained on ImageNet classification. In Tab. 7, we demonstrate that such training strategy can further increase generalization to unseen test features (i.e., DeiT (ViT-B)), increasing overall performance across test feature extractors. We note that this strategy comes at the price of a small degradation of quality on DINOv2 features when compared against the encoder-specific upsampler, which we attribute to overfitting on the encoder-specific positional encoding artifacts of the latter.
Table 7: Multi-Backbone Training. Training on multiple feature extractors achieves comparable performance on seen backbones while improving generalization to unseen ones, outperforming encoder-specific training in cross-backbone evaluation.
DINOv2 SigLIP DeiT Average
AnyUp trained on:
mIoU
($\uparrow$)
Acc
($\uparrow$)
RMSE
(abs) ($\downarrow$)
RMSE
(rel) ($\downarrow$)
mIoU
($\uparrow$)
Acc
($\uparrow$)
RMSE
(abs) ($\downarrow$)
RMSE
(rel) ($\downarrow$)
mIoU
($\uparrow$)
Acc
($\uparrow$)
RMSE
(abs) ($\downarrow$)
RMSE
(rel) ($\downarrow$)
mIoU
($\uparrow$)
Acc
($\uparrow$)
RMSE
(abs) ($\downarrow$)
RMSE
(rel) ($\downarrow$)
DINOv2 62.16 81.37 0.4755 0.3378 58.51 78.36 0.9111 0.5780 53.93 75.90 0.6467 0.497 58.20 78.54 0.6778 0.4709
Multi Backbone 62.04 81.30 0.4767 0.3392 59.52 79.07 0.9088 0.5776 54.83 76.56 0.6418 0.4886 58.80 78.98 0.6758 0.4685
@@ 5.3 Ablation Study
Table 9: Impact of Filter Basis Size on Performance. Scores highlighted from best (gray) to worst (white). We show the previous state-of-the-art results in a separated line for comparison.
$M$ mIoU ($\uparrow$) Acc ($\uparrow$) RMSE (abs) ($\downarrow$) RMSE (rel) ($\downarrow$)
256 62.19 81.38 0.4765 0.3371
128 62.16 81.37 0.4755 0.3378
64 62.09 81.34 0.4759 0.3380
32 62.07 81.31 0.4779 0.3402
16 62.09 81.32 0.4773 0.3404
8 62.00 81.27 0.4792 0.3415
0 61.97 81.23 0.4791 0.3441
SOTA 62.15 81.32 0.4816 0.3413
To further explore this property and quantify the effectiveness of our proposed feature-agnostic layer (Sec. 4.1), we vary the numbers of basis filters in an additional experiment. As can be seen in Tab. 9, upsampling performance increases with larger filter basis size in the feature-agnostic layer, thus validating our hypothesis that a successful upsampler also needs to be able to see the low resolution feature map to distinguish between image areas that are of similar color but of different semantics. We decide to use a filter basis size of $M=128$ in our implementation.
@@ 6 Limitations
FeatSharp (Ranzinger et al., 2025) combines feature upsampling with explicitly de-noising the input features from positional encoding artifacts (Yang et al., 2024). We note that this learned per-model denoising step on the input features can be easily prepended to our pipeline but as the learned denoising weights have not been made public at the time of submission, we do not include this analysis in our experiments. Further, through our training strategy on image parts, our model is effectively also trained to mostly ignore such positional encoding artifacts.
@@ 7 Conclusion
We introduced AnyUp, a first-of-its-kind method for feature upsampling from any resolution to any resolution, which generalizes to feature representations that it was not trained on. Key technical novelties include a feature-agnostic layer, windowed attention, and a training strategy, which work together to achieve state-of-the-art upsampling quality. We make our code and models publicly available at https://github.com/wimmerth/anyup.
@@ Appendix A Additional Details on Related Work
Appendix B Additional Experiments
B.1 Performance Analysis
We summarize performance in terms of runtime and peak memory usage in Tab. 10 (measured on a single NVIDIA A100 GPU). Global attention requires $H\times W\times h\times w$ computations, while window attention, using windows sized $\sigma$ (relative to the input size), significantly reduces the requirement to only $H\times W\times\sigma h\times\sigma w$ computations. In practice, we base our implementation on the $\mathcal{N}$ATTEN library (Hassani et al., 2025), which implements CUDA kernels for several architectures that make use of the sparsity in window attention and set $\sigma=0.2$. We further extend our code to enable trade-off between memory and speed through chunked computation of the attention that enables even larger input images and feature maps.
Table 10: Performance Analysis. AnyUp benefits from the sparsity of the window attention.
(a) Parameter count and inference time (ms) / FLOPs comparison (GFLOPs) by resolution.
Model # Params (M) $112^{2}$ $224^{2}$ $448^{2}$
Time (ms) FLOPs Time (ms) FLOPs Time (ms) FLOPs
FeatUp 0.2 8.6 2.6 20.6 10.4 73.7 41.6
LoftUp 4.3 8.2 47.1 33.5 204.1 201.2 1065.3
JAFAR 0.7 4.3 5.5 19.6 26.8 113.9 186.1
AnyUp 0.8 2.7 5.2 12.8 20.6 92.4 82.4
(b) Memory usage (GB) by resolution for forward / backward passes.
Model $112^{2}$ $224^{2}$ $448^{2}$
Fwd Bwd Fwd Bwd Fwd Bwd
FeatUp 0.3 0.4 0.9 1.4 3.4 5.6
LoftUp 0.2 0.7 0.8 2.8 7.9 21.1
JAFAR 0.2 0.5 0.6 2.7 6.9 22.3
AnyUp 0.2 0.8 0.8 3.3 3.3 12.9
B.2 Comparison against FeatUp’s Implicit Upsampling
Table 11: Comparison against FeatUp (implicit). Feature space preservation comparison with pre-trained linear probes for DINOv2 ViT-S features when upsampling to 224px.
Semantic Segmentation Depth Estimation
Bilinear 35.54 70.20 0.510 0.818
Guided Filter 32.17 68.25 0.537 0.802
FeatUp 35.60 71.40 0.510 0.816
   $\rightarrow$ implicit 36.70 70.95 0.513 0.818
LoftUp 3.82 44.35 0.534 0.803
JAFAR 37.52 72.64 0.505 0.822
AnyUp 37.91 72.78 0.502 0.822
Due to the computational burden of FeatUp’s proposed implicit upsampling, where high-resolution features are optimized in a NeRF-like manner at inference time, it is infeasible to train linear probes for semantic segmentation or depth/normal estimation, as done in Sec. 5.1. Instead, we analyze its performance when used with pre-trained linear probes from DINOv2, as in Tab. 5. We show the results in Tab. 11. Note, that we evaluate on images of 224px instead of 448px used in Tab. 5, as the inference-time optimization of FeatUp is too costly to run on higher resolutions. As in Tab. 5, AnyUp outperforms all prior works when using pre-trained probes.
Appendix C Implementation Details
We train our method on the ImageNet dataset (Krizhevsky et al., 2017) for a total of 100,000 training steps. We choose a batch size of 4 with 4 random local crops per training image. Training takes around 5 hours on a single NVIDIA-H100 GPU. We use the AdamW optimizer (Loshchilov and Hutter, 2019) with a learning rate of 2e-4 and a batch size of 4.